# Road Graph Generator: Mapping roads at construction sites from GPS data

### v2

@@ Appendix A Nomenclature
Notion
Explanation
GPS
Global positioning system.
Intersection
A point in space where at least three roads meet.
Trip
A series of consecutive GPS points from loading a truck until a new trip starts (the truck is re-loaded).
$(lat,lon)$
Coordinates expressed as latitude and longitude.
$(x,y)$
Coordinates expressed in meters with the origin set to the minimal values of latitude and longitude in the dataset.
DTW
Dynamic time warping.
@@ Appendix B Algorithm and data processing parameters
Symbol
Description
Default
Pre-processing parameters
$D_{\mathrm{endpoints}}$
@@
Distance driven removed at the end of each trip.
100 m
$d_{\mathrm{spline}}$
@@
Spline degree used for interpolation.
1
$\Delta D_{\mathrm{interp}}$
Resolution applied during interpolation.
10 m
$N_{\mathrm{min}}$
@@
Minimum number of data points required to define a trip.
4
$\Delta t_{\mathrm{gap}}$
Time threshold (in minutes) for splitting a trip into two separate trips if the vehicle is idle.
5 min
$\Delta\phi_{\mathrm{gap}}$
@@
Angular threshold for splitting a trip into two separate trips based on directional changes.
${}^{\circ}$
Intersection parameters
$N_{\mathrm{res}}$
Resolution of grid cells (height and width).
5 m
$D_{\mathrm{nbr}}$
@@
Maximum distance between the centers of grid cells to consider them neighboring.
20 m
$\Delta\phi_{\mathrm{thr}}$
Threshold for angle variation within neighboring cells to identify intersection candidates.
${}^{\circ}$
$D_{\mathrm{int\_clust}}$
@@
Clustering distance for merging intersection candidates.
15 m
L
@@
Width of extremity annulus used for clustering.
25 m
R
Inner radius of extremity annulus used for clustering.
30 m
$D_{\mathrm{passing}}$
@@
Maximum distance allowed between the trip and the intersection center to include the GPS points belonging to that trip when validating an intersection.
15 m
$D_{\mathrm{ext\_clust}}$
@@
Clustering distance for points in extremity annulus.
30 m
$N_{\mathrm{ext\_clust}}$
Minimum number of points in a cluster.
5
Loading and drop-off nodes parameters
$\epsilon_{\mathrm{load/dump}}$
Radius for expanding clusters in DBSCAN when merging loading or drop-off locations.
0.001
$D_{\mathrm{load/dump}}$
@@
Distance threshold for merging loading or drop-off locations.
100 m
Edge parameters
$D_{\mathrm{node}}$
@@
Maximum distance between GPS track and the node center to consider the trip as passing through that node.
30 m
$\epsilon_{\mathrm{road}}$
@@
Radius for expanding clusters in DBSCAN in road segment detection.
12 m
$N_{\mathrm{min\_road}}$
@@
Minimal sample count in DBSCAN to form a cluster in road segment detection.
5
Table 3: Algorithm and data-processing parameters.

### v3

@@ Appendix A Nomenclature
Notion Explanation
GPS Global positioning system.
Intersection A geographical point where at least three roads meet.
Trip A sequence of consecutive GPS data points representing the movement of a truck from its loading point until it is reloaded, signifying a complete transportation cycle.
$(lat,lon)$ Coordinates expressed as latitude and longitude.
$(x,y)$ Coordinates expressed in meters with the origin set to the minimal values of latitude and longitude in the dataset.
DTW Dynamic time warping.
@@ Appendix B Algorithm and data processing parameters
Data preprocessing parameters
Parameter Description and tuning guidance
@@
$D_{\mathrm{endpoints}}$ Distance driven removed at the end of each trip to reduce noise.
Default: 100 m.
Tuning: Increase to discard noisy points at trip endpoints, especially if GPS accuracy decreases at low speed. Decrease to retain more data near loading/unloading points.
@@
$d_{\mathrm{spline}}$ Spline degree used for interpolation. Higher degrees may lead to overfitting and inaccuracies.
Default: 1 (linear interpolation).
Tuning: Increase for smoother paths, but be cautious of overfitting to noise. Use 1 for reliable, simple interpolation.
$\Delta D_{\mathrm{interp}}$ Resolution for interpolating GPS paths.
Default: 10 m.
Tuning: Increase to reduce the computational cost and memory. Keep or decrease to retain detail in the paths.
@@
$N_{\mathrm{min}}$ Minimum number of data points required to define a valid trip.
Default: 10 points.
Tuning: Increase to exclude very short trips irrelevant for road inference. Decrease to capture brief trips if necessary.
$\Delta t_{\mathrm{gap}}$ Time threshold to split trips if the vehicle is idle for too long.
Default: 5 min.
Tuning: Increase if frequent long pauses should count as part of the trip. Decrease to reflect shorter idles as separate trips.
$\Delta d_{\mathrm{gap}}$ Distance threshold to split trips if consecutive GPS updates are too far apart, indicating a noisy signal.
Default: 2000 m.
Tuning: Increase for large distances with sparse updates that need connecting. Decrease for noisy updates causing errors in road segments.
@@
$\Delta\phi_{\mathrm{gap}}$ Angular threshold to split trips based on directional changes.
Default: 0.
Tuning: Increase if significant directional changes accurately represent the road layout. Decrease to split trips at sharp turns (e.g., U-turns) for better road inference.
Table 3: Parameters used in data preprocessing.
2D histograms of heading directions
Parameter Description and tuning guidance
$N_{\mathrm{res}}$ Grid cell resolution for the 2D histogram construction of heading directions.
Default: 5 m.
Tuning: Increase to speed up processing by reducing the cell count. Decrease for finer granularity in detecting directional changes.
Table 4: Parameters used in 2D histograms of heading directions.
Identification of candidate intersections
Parameter Description and tuning guidance
@@
$D_{\mathrm{nbr}}$ Maximum distance between grid cell centers to consider them as neighbors. Neighboring cells are used to assess directional dissimilarity between the intersection center and its neighbors based on the principal heading directions.
Default: 20 m.
Tuning: Increase to consider larger intersections (and vice versa).
$\Delta\phi_{\mathrm{thr}}$ Threshold for the angular difference in principal heading direction between neighboring cells to identify them as intersection candidates.
Default: 1.4.
Tuning: Decrease to detect smaller angular changes indicating intersections. Increase to focus on significant direction changes.
@@
$D_{\mathrm{int\_clust}}$ Clustering distance for merging nearby intersection candidates.
Default: 15 m.
Tuning: Increase to merge more distant intersection candidates. Decrease to separate closely located ones.
Table 5: Parameters used for identification of intersection candidates.
Validation of intersections
Parameter Description and tuning guidance
$R$ Inner radius of the extremity annulus for validating intersections.
Default: 30 m, 100 m.
Tuning: Increase to capture intersections with wider centers. Decrease to detect more compact intersections and account for smaller roads leaving the intersection.
@@
$L$ Width of the extremity annulus used in clustering, with the outer radius defined as $R+L$.
Default: 25 m.
Tuning: Increase to capture more GPS points around the intersection. Decrease to avoid merging roads that connect further away into a single cluster.
$N_{\mathrm{max\_val}}$ Maximum number of GPS points in the extremity annulus used for each intersection validation. .
Default: 1000.
Tuning: Increase for potential increase in accuracy. Decrease for faster performance.
@@
$D_{\mathrm{passing}}$ Maximum distance between the trip and the intersection center for including GPS points in intersection validation.
Default: 15 m.
Tuning: Increase for wide intersections to capture all trips passing through. Decrease to exclude trips passing nearby but not through the intersection.
@@
$D_{\mathrm{ext\_clust}}$ Distance for clustering GPS points within the extremity annulus to represent outgoing roads.
Default: 30 m.
Tuning: Increase to merge distant or scattered GPS points as a single road. Decrease to prevent merging multiple roads into one.
$N_{\mathrm{ext\_clust}}$ Minimum number of points required in a cluster within the extremity annulus to consider it as a road passing through the intersection.
Default: 5 points.
Tuning: Increase for stricter road validation, enhancing the reliability of the algorithm. Decrease to recognize clusters with fewer points and include less frequently traveled roads.
Table 6: Parameters used for validation of intersections.
Loading and drop-off nodes
Parameter Description and tuning guidance
@@
$D_{\mathrm{load/dump}}$ Distance threshold for merging loading or drop-off locations.
Default: 100 m.
Tuning: Increase to combine more distant locations, resulting in larger aggregated zones. Decrease to identify more separated locations.
Table 7: Parameters used for identification of loading and drop-off nodes.
Road inference
Parameter Description and tuning guidance
@@
$D_{\mathrm{node}}$ Maximum distance between a GPS track and a node center for the trip to be considered as passing through that node (intersection or a loading/drop-off point).
Default: 30 m.
Tuning: Increase to include roads passing further from the node center, such as at wide intersections with dispersed tracks. Decrease to exclude roads that are only nearby.
@@
$\epsilon_{\mathrm{road}}$ Radius for DBSCAN clustering during road segment detection.
Default: 15 m.
Tuning: Increase to merge more spread-out GPS points, possibly connecting separate road segments. Decrease for tighter clustering, focusing on clearer road definitions.
@@
$N_{\mathrm{min\_road}}$ Minimum sample count for DBSCAN to form a road segment cluster.
Default: 5 points.
Tuning: Increase to require more GPS points for defining a road, which can eliminate noise but may miss roads with sparse data. Decrease to detect more roads, even those with less data coverage.
Table 8: Parameters used for road inference.